OpenStudy (anonymous):

.80+.40=12.8 c+t=24 what is c

OpenStudy (anonymous):

what do I do from here @pinkbubbles

OpenStudy (anonymous):

here's the word problem Alex stocks up for winter. He buys 24 cans of vegetables. He pays 80 cents per can for tomatos and 40 cents per can for corn, for a total cost of $12.8

OpenStudy (anonymous):

\[\text{Solve}[\{0.8 c+0.4 t=12.8,c+t=24\}] \]Work it out. c should equal 8 .

OpenStudy (anonymous):

but how do I work it out

OpenStudy (anonymous):

c + t =24, thus, t = 24 - c Plug 24-c for each occurrance of t in the first equation and solve the result for c.

OpenStudy (anonymous):

so then it would be 3?

OpenStudy (anonymous):

Sorry, occurrance should have been spelled, occurrence.

OpenStudy (anonymous):

so u said that c would equal 8 then wouldn't t equal 3

OpenStudy (anonymous):

0.4 (24 - c) + 0.8 c = 12.8 0.8 c - 0.4 c + 9.6 =12.8

OpenStudy (anonymous):

so what do I do next

OpenStudy (anonymous):

0.4 c = 12.8 - 9.6 0.4 c = 3.2 c = 8

OpenStudy (anonymous):

so how did u get 9.6

OpenStudy (anonymous):

so is 8 the answer or the value of c

OpenStudy (anonymous):

0.4 (24 - c) + 0.8 c = 12.8 9.6 + 0.8 c = 12.8

OpenStudy (anonymous):

"so is 8 the answer or the value of c ? " Yes.

OpenStudy (anonymous):

so the answer would be 16 right